# Affordance-based Robot Manipulation with Flow Matching

### v4

@@ IV-B1 Baseline Studies
IV-B2 Main Results
Table II presents the results of flow matching policy on our ADLs testing dataset for robot trajectory learning, comparing against baselines. We use two metrics for evaluation: (i) the error of 2D and 3D trajectory estimation, and (ii) the inference time, performed with RTX 4090 GPU. The trajectory error is averaged on each point of the trajectory.
Four observations could be made from this result:
  • Generation Quality: Flow matching (CNN-based, 16 steps) outperforms diffusion policy and Transformer baselines in terms of 2D and 3D trajectory prediction accuracy. The Transformer behavior cloning achieves marginal precision. This is expected as it is hindered by the nature of multi-modal action distribution, causing the averaging out across non-convex spaces.
  • Inference Time: We have two observations here: i) Table II showcases that flow matching with 16 steps achieves faster inference time compared to diffusion policy with 16 steps. We hypothesize that flow matching with linear pointwise flows generates straighter flows than DDPM and DDIM, and thus causes faster inference. ii) More importantly, Table II also showcases that 1-step flow matching (error: $1.031$cm, time: $13.228$ms) has achieved comparable performance as 16-step DDIM (error: $2.266$cm, time: $98.680$ms), but noticeably lowered inference time roughly by 85%. We hypothesize that this is because diffusion models solve a stochastic differential equation with a series of discrete steps to progressively refine the generated sample. Contrarily, flow matching trains continuously normalize flow models, leading to no significant improvements when increasing inference steps. Thus 2-step flow matching has achieved comparable performance as 16-step diffusion policy, which considerably reduces the inference time for closed-loop robot manipulation. This is in line with some results obtained in the image generation domain. As pointed out by Stable Diffusion 3 [6], flow matching performs better with fewer inference steps than diffusion policy, especially when faster inference is required. The extensive evaluation on multiple robot benchmarks in the next section will further reinforce our argument.
IV-B3 Real-World Robot Experiments
We deploy flow matching, DDPM and Transformer policies on real robot manipulation for evaluation. We carry out 50 replications of trials for each baseline. We use a KINOVA Gen3 arm and an Azure Kinect camera for real-world robot experiments. Details can be found in the supplementary video.
From Table III, we can see that flow matching outperforms other baselines. We also observe that most failures occurred due to some out-of-distribution factors, including background and significant camera view changes. Larger data size and more diversity should alleviate this problem.
RGB images with/without affordances: We also investigate how the affordance would guide the flow matching policy. We trained flow matching taking the raw RGB images and language tokens as input. This is similar to our proposed method but without the intermediate stage of affordance learning. It also involves some resemblance while not being entirely identical to the method in [2]. Interestingly, comprehensive examinations reinforce the argument that flow matching could handle multimodal action distribution. Fig. 7 shows one example. From the left figure, we can see that when training a policy without affordances, the predicted trajectory (yellow) of moving the towel toward the trash for sweeping could be detached from the ground truth (red), but still a reasonable solution that allows for a successful robot execution. With affordance guidance, the prediction is closely aligned with the truth (Fig. 7-right). We also observe that for applications that demand higher precision in manipulation, like grasping the toothbrush handle, affordances offer greater guidance for shaping the manipulation policy.
Figure 7: Ablations of using RGB images with/without affordances for policy training. Visual affordance guides the flow matching policy to generate a trajectory closely aligned with the truth. The policy without affordances might generate a trajectory detached from the ground truth, but potentially still a reasonable solution. This reinforces the argument that flow matching policy could handle multimodal robot action distributions.
IV-B4 Ablations
@@
Trajectory Reprsention: We empirically test trajectory representation with 8, 16, 32 and 64 waypoints. More waypoints are not necessary, while fewer waypoints are unable to entirely encapsulate the complete long-horizon trajectories. We have found that, in general, the trajectory representation does not wield a significant influence on flow matching performance. The performance reported in the above main results section is achieved by using 32 waypoints.
@@
(Inference Step) ($16$-step) $\uparrow$ ($16$-step) $\uparrow$ BC $\uparrow$ RGB $\uparrow$
@@
IV-C Comparisons between flow matching and diffusion policy
To further investigate the performance of flow matching compared to diffusion policy, we benchmark the proposed methods on three more datasets which include closed-loop 6D robot actions and gripper actions: (i) Push-T [7], (ii) Franka Kitchen [14], and (iii) Robomimic [33].
  • Franka Kitchen contains 7 objects for interaction and comes with a human demonstration dataset of $566$ demonstrations, each completing 4 tasks in arbitrary order. The goal is to execute as many demonstrated tasks as possible, regardless of order. The training takes state-based inputs, and outputs closed-loop robot joint actions and gripper actions.
  • Robomimic consists of 5 tasks with a proficient human teleoperated demonstration dataset. We specifically focus on the transport task which includes $200$ demonstrations. The policy takes state-based inputs, and outputs closed-loop robot joint actions and gripper actions.
For each benchmark, the evaluation has been carried out across $500$ different environment initial conditions, using the last checkpoint of each policy with $10$ trials of replications. Thus, $5,000$ trials have been carried out in total per policy and benchmark. Variation is added on random initial conditions for the robot and object states. We respectively report the best and average performance in the $10$ trials of replications of the last checkpoint. All state-based tasks are trained for $4,500$ epochs, and image-based tasks for $3,000$ epochs. Table V shows the hyperparameters we have used in flow matching and diffusion policy. Table VI shows the task summary.
  • Generation Quality: Flow matching outperforms DDPM in all three benchmarks. The performances of flow matching and DDIM are comparable, where flow matching performs marginally better in most cases.
  • Inference Time: Fig. 8 shows how the number of inference steps affects the performance of flow matching and diffusion policy. We can observe that diffusion policy showcases better performances when applying more inference iterations with a trade-off of longer inference time, as it requires a series of discrete steps to progressively refine the generated sample. Contrarily, flow matching has not shown significant improvements when increasing inference steps. Therefore, flow matching considerably reduces the inference time for closed-loop robot manipulation. In the Push-T benchmark in Fig. 8, 2-step flow matching (coverage: $0.8803$, time: $13.098$ms) has achieved comparable performance as 16-step diffusion policy with DDIM (coverage: $0.8801$, time: $98.268$ms), but noticeably lower inference time roughly by 86%.
Figure 8: Generation quality and inference time comparison of flow matching and diffusion policy for varying values of inference steps.
@@ V Limitations
The generation quality of flow matching and diffusion policy for robot manipulation are generally comparable. Although we see only a marginal improvement in flow matching in most cases, we would like to highlight that the focus of this work is not to outperform state-of-the-art general robot manipulation research. Instead, we have systematically studied the flow matching framework, which provides an alternative to diffusion policies for robot manipulation. We can not overlook the additional advantages of flow matching, including stable training, easy implementation, and most importantly, significantly better performance and faster inference with fewer inference steps than diffusion policy, suggesting forsaking the stochastic construction of diffusion policy in favor of learning the probability path more directly as in flow matching.
@@
We have evaluated how varying inference steps affect the performance of flow matching and diffusion policy. The original diffusion policy research [3] adopts a large $100$ training and inference steps with DDPM in their experiments. Based on our evaluation in Fig. 8 and results from other research [4], we can observe that beyond $8$ steps, further increasing steps have only a marginal impact on the performance of diffusion policy, but with a trade-off of significantly longer inference time. We are aware of recent competitive research on one-step diffusion policy with distillation [38] and shortcut models [8]. In this paper, we primarily focus on conducting a comparative analysis of the fundamental architectures underlying flow matching and diffusion policy with DDPM and DDIM.
We have not explored applying flow matching for robot action generation in the reinforcement learning setting. Recent research [51] has modeled a policy as a return-conditional flow matching model with classifier-free guidance to eliminate many of the complexities that come with traditional offline reinforcement learning and outperformed standard imitation learning. These methods are anticipated to be seamlessly adaptable to robotic manipulation scenarios; however, comprehensive validation through practical simulations and real-world experimentation remains imperative.

### v5

@@ IV-B2 Main Results
IV-B3 Ablations
@@
Figure 7: Ablations of using RGB images with/without affordances for policy training. Visual affordance guides the flow matching policy to generate a trajectory closely aligned with the truth. The policy without affordances might generate a trajectory detached from the ground truth, but still a reasonable solution. This reinforces the argument that flow matching policy could handle multimodal robot action distributions.
@@
Trajectory Representation: We empirically test trajectory representation with 8, 16, 32 and 64 waypoints. More waypoints are not necessary, while fewer waypoints are unable to entirely encapsulate the complete long-horizon trajectories. We have found that, in general, the trajectory representation does not wield a significant influence on flow matching performance. The performance reported in the above main results section is achieved by using 32 waypoints.
IV-C Real-World Robot Evaluation
We deploy our flow matching policy, DDPM and Transformer policies on real robot manipulation for evaluation. We carry out 50 replications of trials for each baseline. We use a KINOVA Gen3 arm and an Azure Kinect camera for real-world robot experiments. Details can be found in the supplementary video. Besides, we have respectively compared the proposed parameter-efficient affordance learning and flow matching policy against SoA baselines. In this section, we also investigate whether our framework can seamlessly unify affordance learning and action generation.
Affordance-based policy: From Table III, we can see that flow matching outperforms DDPM and Transformer baselines.
End-to-end learning (RGB images without affordances): We investigate how the affordance would guide the flow matching policy. We trained flow matching taking the raw RGB images and language tokens as input. This is similar to our proposed method but without the intermediate stage of affordance learning. It also involves some resemblance while not being entirely identical to the method in [black2024pi_0] in an end-to-end learning manner. Interestingly, comprehensive examinations reinforce the argument that flow matching could handle multimodal action distribution. Fig. 7 shows one example. From the left figure, we can see that when training a policy without affordances, the predicted trajectory (yellow) of moving the towel toward the trash for sweeping could be detached from the ground truth (red), but still a reasonable solution that allows for a successful robot execution. With affordance guidance, the prediction is closely aligned with the truth (Fig. 7-right, Table). We also observe that for applications that demand higher precision in manipulation, like grasping the toothbrush handle, affordances offer greater guidance for shaping the manipulation policy.
Off-the-shelf VLM-based trajectory generation: Our method uses an extra low-level module (flow matching) to learn a robot policy built on the high-level VLM reasoning. We have also conducted preliminary research to test another popular line of research, which uses a VLM to generate robot actions directly, as customary in the literature [huang2023voxposer, tanneberg2024help]. No constant performance was achieved without carefully designed prompts or finetuning. This result is expected as several recent state-of-the-art research (e.g., HAMSTER [li2025hamster], $\pi$[black2024pi_0]) have shown that VLA models with robot learning modules outperform VLM-based trajectory generation.
@@ IV-C Real-World Robot Evaluation
(Inference Step) ($16$-step) $\uparrow$ ($16$-step) $\uparrow$ BC $\uparrow$ end-to-end $\uparrow$
@@
IV-D Comparisons between flow matching and diffusion policy
To further investigate the performance of flow matching compared to diffusion policy, we benchmark the proposed methods on three more datasets which include closed-loop 6D robot actions and gripper actions: (i) Push-T [florence2022implicit], (ii) Franka Kitchen [gupta2019relay], and (iii) Robomimic [robomimic2021].
  • Franka Kitchen contains 7 objects for interaction and comes with a human demonstration dataset of $566$ demonstrations, each completing 4 tasks in arbitrary order. The goal is to execute as many demonstrated tasks as possible, regardless of order. The training takes state-based inputs and outputs closed-loop robot joint actions and gripper actions.
  • Robomimic consists of 5 tasks with a proficient human teleoperated demonstration dataset. We specifically focus on the transport task which includes $200$ demonstrations. The policy takes state-based inputs and outputs closed-loop robot joint actions and gripper actions.
For each benchmark, the evaluation has been carried out across $500$ different environment initial conditions, using the last checkpoint of each policy with $10$ trials of replications. Thus, $5,000$ trials have been carried out in total per policy and benchmark. Variation is added on random initial conditions for the robot and object states. We respectively report the best and average performance in the $10$ trials of replications of the last checkpoint. All state-based tasks are trained for $4,500$ epochs, and image-based tasks for $3,000$ epochs.
[chi2023diffusion] has reported that setting prediction and action horizons greater than 1 helps the policy generate more consistent actions and compensate for idle portions of demonstrations; however, excessively long horizons can degrade performance due to slower reaction times. To ensure fair comparisons, we have selected the same corresponding hyperparameters across the flow matching policy and baseline methods to remain consistent. Table V shows the hyperparameters we have used in flow matching and diffusion policy. Table VI shows the task summary.
  • Generation quality: The performances of flow matching and DDPM, DDIM and the score-based model are comparable, where flow matching performs marginally better in most cases. We observe that score-based models can achieve higher peak performance, but their average performance tends to be less consistent.
  • Inference time: Fig. 8 shows how the number of inference steps affects the performance of flow matching and diffusion policy. We can observe that diffusion policy showcases better performances when applying more inference iterations with a trade-off of longer inference time, as it requires a series of discrete steps to progressively refine the generated sample (solid line). Score-based models with black-box numerical ODE samplers, as in the original research paper [song2020score], need adaptive steps to achieve optimal results (around 100300100-300). Contrarily, flow matching has not shown significant improvements when increasing inference steps. From 8 steps onward, the performance of both flow matching and the diffusion policy increases only slightly with additional steps. However, the inference time (dotted line, we use frequency here) increases proportionally with the number of inference steps (dotted line). Therefore, flow matching considerably reduces the inference time for closed-loop robot manipulation. In the Push-T benchmark in Fig. 8, 2-step flow matching (coverage: $0.8803$, time: 13.09813.098ms) has achieved comparable performance as 16-step diffusion policy with DDIM (coverage: $0.8801$, time: 98.26898.268ms), but noticeably lower inference time roughly by 86%.
Based on the above experimental observations, we conclude that flow matching formulates the robot policy generation problem as learning a deterministic or near‐deterministic vector field (via an ordinary differential equation) that maps noise (or a simple prior) to target action distributions—this ODE-based framing avoids iterative stochastic denoising chains (as in diffusion policy) and yields fast, single- or few-step inference, which is crucial in closed-loop robotic settings. Besides, in manipulation tasks, the target action distribution is often multimodal (e.g., multiple possible grasp poses, multiple contact sequences) but still has strong structure (e.g., physics constraints, smooth trajectories through state‐action space). Flow matching’s vector‐field representation permits modeling such multimodality while preserving smooth interpolation and strong inductive bias towards consistent flows—thus potentially better capturing the structured trajectory manifold than, say, a GAN which may collapse modes, or a diffusion model which requires many steps. Moreover, because robotic manipulation is extremely sensitive to inference latency, stability, and feedback responsiveness, flow matching’s fewer-step generation is advantageous. The explicit mapping of a vector field lends itself more naturally to feedback correction and incorporation of geometric/physical structure (e.g., SE(3) equivariance) than purely stochastic models. Overall, flow matching combines fast inference and structured multimodal representation, making it a compelling choice for manipulation domains.
During our experiments, we observed that most failures were caused by out-of-distribution factors. In real-world experiments, variations such as background changes or significant camera view shifts often led to failures of the 3D flow-matching policy. Increasing both the dataset size and diversity could help mitigate this issue. Unlike the simulation benchmarks, we performed open-loop manipulation in the real world, meaning that a failed first grasp attempt was counted as a failure; closed-loop manipulation could potentially address this limitation. We also found that tasks requiring higher manipulation precision, such as grasping a toothbrush handle, exhibited a higher probability of failure. In simulation benchmarks, out-of-distribution failures were also observed. For example, in the Push-T benchmark, only successful trajectories were included during training, so the policy struggled with edge cases, such as a T-block stuck in a corner. Extending the imitation policy to an online or offline reinforcement learning framework may help address these limitations.
@@ V Limitations
In this section, we provide a detailed analysis of the advantages and limitations of our proposed affordance-based flow matching approach.
@@
Inference time: The generation quality of flow matching and diffusion policy for robot manipulation are generally comparable. Although we see only a marginal improvement in flow matching in most cases, we would like to highlight that the focus of this work is not to outperform state-of-the-art general robot manipulation research. Instead, we have systematically studied the flow matching framework, which provides an alternative to diffusion policies for robot manipulation. We can not overlook the additional advantages of flow matching, including stable training, easy implementation, and most importantly, significantly better performance and faster inference with fewer inference steps than diffusion policy, suggesting forsaking the stochastic construction of diffusion policy in favor of learning the probability path more directly as in flow matching.
@@
We have evaluated how varying inference steps affect the performance of flow matching and diffusion policy. The original diffusion policy research [chi2023diffusion] uses a large $100$ training and inference steps with DDPM in their experiments. Based on our evaluation in Fig. 8 and results from other research [chisari2024learning], we can observe that beyond $8$ steps, further increasing steps have only a marginal impact on the performance of diffusion policy, but with a trade-off of significantly longer inference time. We are aware of recent research on one-step diffusion policy with distillation [prasad2024consistency] and shortcut models [frans2024one]. We primarily focus on conducting a comparative analysis of the fundamental architectures underlying flow matching and vanilla diffusion policy with DDPM, DDIM and the score-based model.
Figure 8: Generation quality (solid line) and inference time (dotted line, we use frequency here) comparison of flow matching and diffusion policy for varying values of inference steps.
Deterministic and stochastic tasks: Flow Matching is inherently more deterministic than diffusion-based approaches, which can have implications in highly stochastic tasks—i.e., tasks where the same actions can lead to different outcomes due to environmental randomness or uncertainty, such as manipulating deformable objects, pushing objects on uneven surfaces, or navigating in dynamic environments. The impact of flow matching’s determinism also depends on the training dataset: with a diverse dataset covering a wide range of outcomes, flow matching can learn to predict the most likely trajectories effectively, whereas in limited or biased datasets, flow matching may fail to capture rare but plausible outcomes.
Trajectory smoothness: Although our flow matching framework effectively learns Cartesian-space or joint-space trajectory representations, it does not explicitly enforce second-order smoothness or torque feasibility during training or inference. In practice, the trajectories are executed through standard robot APIs (e.g., Kinova), which apply spline-based interpolation and internal velocity/acceleration limits to ensure dynamically feasible and smooth motion. Nevertheless, incorporating explicit smoothness or dynamic-feasibility constraints—such as velocity or acceleration regularization terms, or post-processing filters—could further improve motion consistency and physical plausibility. However, as noted in recent works [romer2024diffusion, zeng2025optimization], these additions typically increase computational complexity, highlighting a trade-off between control fidelity and real-time efficiency. Exploring this direction is a valuable avenue for future research.
Transportation cost: While our method successfully leverages basic flow matching to generate robust and feasible affordance-based trajectories with a high success rate, we acknowledge that it does not explicitly pursue the minimal geometric transportation cost. More advanced flow techniques, such as Rectified Flow Matching (RFM) [liu2022rectified] or methods based on Optimal Transport (OT) [villani2021topics], are mathematically designed to produce straight-line mappings that minimize this cost. However, implementing these methods introduces significant training and preprocessing overhead, often necessitating complex iterative or global optimization schemes. Given that our primary goal is to explore the fundamental structural efficacy and computational tractability and use affordances to shape the target distribution, we adopt the more efficient basic flow matching objective. We are deferring the complex, multi-objective optimization that simultaneously constrains geometric factors like path length and crucial practical constraints to future work.
Scalability: The current scope of our work presents several avenues for future generalization. The ADLs dataset was constructed in a controlled setting. We acknowledge this limited variation necessitates future work in demonstrating cross-domain generalization (e.g., varying lighting or camera poses), as acquiring and annotating the vast, diverse datasets required for open-world robustness is often infeasible for typical research settings. Furthermore, while our parameter-efficient prompt tuning leverages the semantic robustness of a pretrained foundation model to learn manipulation affordances, incorporating the spatial and semantic interactions between elements in multi-task scenarios, its ultimate performance against extreme linguistic noise is fundamentally data-driven. We plan to address these challenges in future work by expanding dataset diversity and evaluation rigor. Besides, we are exploring future extensions that incorporate additional modalities (e.g., haptic or force-torque sensing) or dynamics models to provide a richer, dynamics-aware affordance representation.